A pharmacist needs to make a 1% hydrocortisone cream. How many grams of 2.5% hydrocortisone cream should be added to 15 grams of cream base (0% hydrocortisone) to make the 1% cream?
Answer by lwsshak3(11628)

A pharmacist needs to make a 1% hydrocortisone cream. How many grams of 2.5% hydrocortisone cream should be added to 15 grams of cream base (0% hydrocortisone) to make the 1% cream?
..
let x=grams of 2.5% cream to be added
grams of cream base to be added=15
then x+15 = total grams of final 1% cream
.025x+0*15=.01(x+15)
2.5x=x+15
1.5x=15
x=10 grams
ans:10 grams of 2.5% hydrocortisone cream should be added to 15 grams of cream base to make the 1% cream?